OpenStudy (anonymous):

****** Metal ******* use the given trigometric ratio to determine which angle of the triangle is angle A 2.tan A 1 1/3 3. cos A= 0.6 4. cos A = 0.8 5. tan A=0.75 6. sin A = 0.6

Directrix (directrix):

>diagram shown below How far below? :)

OpenStudy (anonymous):

here below

OpenStudy (anonymous):

@johnweldon1993 @jim_thompson5910 hlp

Directrix (directrix):

Is this six different questions or just one question with a lot of given information?

OpenStudy (anonymous):

six different questions @Directrix

Directrix (directrix):

Do you know the right triangle trig ratios?

OpenStudy (anonymous):

yes i do @Directrix

Directrix (directrix):

Look at the attachment. Given: sinA = 8/10 = 4/5. The same is true for the sine of angle 1 in the diagram. So, angle A is angle 1. Remember that the sine of an acute angle in a right triangle is the length of the opposite leg over the length of the hypotenuse.

Directrix (directrix):

@mary_arevalo_boo

OpenStudy (anonymous):

there @Directrix

Directrix (directrix):

tan A = 1 1/3 tan A = 4/3 Tangent is the ratio of sine to cosine. @mary_arevalo_boo What is the sine of angle 1 and what is the cosine of angle 1. Then, divide the sine value by the cosine value and see if it is 4/3?

OpenStudy (anonymous):

sorry i cant do it i forgot @Directrix

OpenStudy (anonymous):

can u show your work please @Directrix

OpenStudy (anonymous):

\[\cos^{-1} \frac{ 1 }{ 3 }\]